OpenStudy (anonymous):

Can someone walk through the steps of solving this? (3m^2n^4)^4

OpenStudy (anonymous):

\[(3m^2n^4)^4\]

OpenStudy (princeharryyy):

3^4.m^8.n^16 is the answer @mkbwashere as power gets multiplied if they are in power while in multiplication they gets added ... hope this helps !!!!

OpenStudy (anonymous):

\[=81m^8n^{16}\]

OpenStudy (anonymous):

\[=(3m^3n^4)^4= (3)^4 \times (m^2)^4 \times (n^4)^4 = 81 \times m^8 \times n^{16} = 81m^8n^{16}\]

OpenStudy (mokeira):

\[3^{4}m ^{8}n ^{16}\]